A 2-kg disk has a radius of 18 cm and rotates with an angular acceleration of 12 rad/s2

about 

an axis through its center and perpendicular to the plane of the disk. Determine the tangential 

force at the rim of the disk.


1
Expert's answer
2021-03-15T20:57:05-0400

Solution:


I = "\frac{1}{2}" MR² = "\frac{1}{2}" (2)(0.18)²

τ = αI  = "\frac{1}{2}" (2)(0.18)²(12)

τ = FR

F = "\dfrac{τ}{R}"


F = "\\dfrac{[\\frac{1}{2} \n2\n1\n\u200b\t\n (2)(0.18)\u00b2(12)]}{0.18}" = 2.16 N